OpenStudy (anonymous):

One More Question: Find the difference quotient of f for this function: f(x) = 1 ---- x + 2 (can you also put a breakdown of how you got the answer? )

OpenStudy (anonymous):

i gues you will differentiate

OpenStudy (anonymous):

?

OpenStudy (dumbcow):

\[=\frac{\frac{1}{(x+h)+2} - \frac{1}{x+2}}{h}\]

OpenStudy (dumbcow):

\[=\frac{\frac{x+2 - (x+h+2)}{(x+2)(x+h+2)}}{h} = \frac{-h}{h(x+2)(x+h+2)}\]

OpenStudy (dumbcow):

cancel the h's and then it is simplified

OpenStudy (anonymous):

follow dumbcows order thaz wat i mean

OpenStudy (dumbcow):

when h=0 , it gives the derivative so mathfrick was right, you just haven't got to that step yet

OpenStudy (anonymous):

do I have to breakdown the denominator and distribut h before i simplify? i can't seem to get the right answer

OpenStudy (dumbcow):

no need to distribute, h/h = 1

OpenStudy (dumbcow):

= -1/(x+2)(x+h+2)

OpenStudy (anonymous):

that's the answer?

OpenStudy (dumbcow):

yes, the answer is just using the def they gave you f(x+h) -f(x)/h all i did was combine fractions and simplified